OpenStudy (anonymous):

Can someone explain how this equation is solved along with the answer? 25^3x+2=125

OpenStudy (anonymous):

Can you double check your equation?

OpenStudy (anonymous):

Is that really 25^3?

OpenStudy (anonymous):

Also, this Equation too. \[\log_{6} x+ \log_{6} (x-5)=2\]

OpenStudy (anonymous):

@zordoloom \[25^{3x+2}=125\]

OpenStudy (anonymous):

Lake log base 25 of both sides you will get 3x+2=3/2 then isolate x you get 3x=-.5 x=-1/6

hartnn (hartnn):

can u write 25 and 125 in powers of 5 ?

OpenStudy (campbell_st):

rewrite the equation in terms of powers of 5 \[(5^2)^{3x + 2} = 5^3\] which becomes \[5^{6x + 4} = 5^3\] just equate the powers now you have the same base and solve for x.

OpenStudy (anonymous):

Take not lake. Sorry

OpenStudy (campbell_st):

for the log question just rewrite it using log laws \[\log(a) + \log(b) = \log (ab)\] so you will have \[\log_{6}(x (x - 5)) = 2\] next raise each term as a power of 6 \[a^{\log_{a}(b)} = b\] so \[6^{\log_{6}(x(x-5)} = 6^2\] or \[x(x -5) = 36\] solve for x